OpenStudy (anonymous):

Find the angle between the given vectors to the nearest tenth of a degree. u = <-5, -4>, v = <-4, -3> @agent0smith

OpenStudy (agent0smith):

a · b = |a| × |b| × cos(θ) | | means magnitude. a*b means dot product

OpenStudy (anonymous):

i'm confused ...

OpenStudy (agent0smith):

Know how to find dot product? It's easy. a = <s, t> and b = <x, y> a·b = s*x + t*y

OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (agent0smith):

We haven't finished the first question... and that one uses the same method.

OpenStudy (agent0smith):

a = <s, t> and b = <x, y> a·b = s*x + t*y do that for u = <-5, -4>, v = <-4, -3> u·v = ?

OpenStudy (anonymous):

i got 1.8

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

How? Multiply the x components together. Multiply the y components together. Add the two together u = <-5, -4>, v = <-4, -3> u·v = (-5)*(-4) + (-4)*(-3)

OpenStudy (anonymous):

i got it as my final answer

OpenStudy (anonymous):

for that 32

OpenStudy (agent0smith):

32 = |u| × |v| × cos(θ) where |u| = sqrt(5^2 + 4^2) |v| = sqrt(4^2 + 3^2) if you did something like this, and found theta you're right. I haven't checked the answer.

OpenStudy (anonymous):

yeah i did that, do you think im right

OpenStudy (agent0smith):

Yes, 1.8 degrees is correct!

OpenStudy (anonymous):

Yay c:

OpenStudy (agent0smith):

u = <6, -2>, v = <8, 24> for this find u·v = ?

OpenStudy (anonymous):

48,-48

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Dot product isn't a vector. you add the numbers.

OpenStudy (anonymous):

Determine whether the vectors u and v are parallel, orthogonal, or neither. u = <6, -2>, v = <8, 24>

OpenStudy (anonymous):

Find the indicated dot product. r = <9, -7, -8>, v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w @agent0smith

OpenStudy (agent0smith):

Same as above. v = <3, 4, 7>, w = <6, -9, 7> v ⋅ w = multiply the x's, y's and z's

OpenStudy (agent0smith):

Same way you did it with the ones above, just this time you have two extra numbers to multiply.

OpenStudy (anonymous):

18,-36,49?

OpenStudy (agent0smith):

Dot product isn't a vector, you add the numbers.

OpenStudy (anonymous):

so 31

OpenStudy (agent0smith):

That's your answer :)

OpenStudy (anonymous):

Express the complex number in trigonometric form. -3 + 3 square root of threei

OpenStudy (agent0smith):

Um, you'll have to find the r, which is the magnitude... sqrt( 3^2 + (3sqrt3)^2 ) and then find the angle. Tan theta = -3/(3sqrt3)

OpenStudy (anonymous):

i got something completly off

OpenStudy (agent0smith):

...?